# The Relative Entropy of Expectation and Price

### v6

@@ III Entropic pricing and hedging
      In the second study, the final prices are modelled as independent binomial variables with possible values $(0,1+c\,dt)$ and probabilities $(w,1-w)$. Examples are plotted for the first underlying with coupon $c_{1}=5\%,10\%,15\%$ and default probability $w_{1}=0.01\%,\ldots,10\%$. The second underlying has coupon $c_{2}=10\%$ and the implied funding rate is plotted as a function of its default probability $w_{2}$. The implied funding rate is capped by the coupons and decreases as the default probability increases.
III.1 Calibration
Calibration generates the implied funding rate $r$ and the price measure $\mathbb{E}^{\phi}$ equivalent to the expectation measure $\mathbb{E}$ with Radon-Nikodym weight:
$\frac{d\mathbb{E}^{\phi}}{d\mathbb{E}\hfill}:=\frac{\displaystyle\exp[-\phi\cdot dq]}{\displaystyle\mathbb{E}[\exp[-\phi\cdot dq]]}$ (46)
where the unit optimal portfolio $\phi$ satisfies the self-funding condition $\phi\cdot q=0$ and the calibration condition:
$0=\mathbb{E}^{\phi}[dq]-qr\,dt$ (47)
All investments are self-funded within the underlying market, and this model does not assume the existence of a risk-free funding rate for investments that are not self-funded. The implied funding rate defined here originates in the optimisation as the Lagrange multiplier for the self-funding condition.
If the underlying market includes a risk-free funding security with initial price $b>0$ and final price $B=b+db$ satisfying $\mathbb{V}[db]=0$, then the calibration condition derives the solution $r=\mathbb{E}[db]/(b\,dt)$ for the implied funding rate. More generally, the implied funding rate is discovered simultaneously with the unit optimal portfolio by iterating the Newton-Raphson scheme:
$\partial r$ $=\frac{\displaystyle(\phi+\mathbb{V}^{\phi}[dq]^{-1}(\mathbb{E}^{\phi}[dq]-qr\,dt))\cdot q}{\displaystyle\mathbb{V}^{\phi}[dq]^{-1}q\,dt\cdot q}$ (48)
$\partial\phi$ $=\mathbb{V}^{\phi}[dq]^{-1}(\mathbb{E}^{\phi}[dq]-q(r+\partial r)\,dt)$
initialised with $r=0$ and $\phi=0$.
@@ III.1 Calibration
As evident in this calibration scheme, the implied funding rate incorporates contributions from all the underlying returns, weighted by their predictabilities so that the least risky returns dominate. This property is demonstrated in the studies of figure 6, which simplify the model with the assumption that the final underlying prices are independent. The calibration conditions decouple to generate a set of independent conditions:
@@
that are solved separately to express the components of the unit optimal portfolio $\phi[r]$ as functions of the implied funding rate $r$. The self-funding condition $\phi[r]\cdot q=0$ is then solved for $r$.
In the first study, the final underlying price $Q_{i}$ is modelled as a normal variable with mean $q_{i}(1+\mu_{i}\,dt)$ and variance $q_{i}^{2}\sigma_{i}^{2}\,dt$, where $\mu_{i}$ is the drift and $\sigma_{i}$ is the volatility. The calibration condition derives the unit optimal portfolio:
@@
and the self-funding condition then derives the implied funding rate:
@@
In the second study, the final underlying price $Q_{i}$ is modelled as a binomial variable with possible values $(0,1+c_{i}\,dt)$ occurring with probabilities $(w_{i},1-w_{i})$, where $c_{i}$ is the coupon and $w_{i}$ is the default probability. The calibration condition derives the unit optimal portfolio:
@@
III.2 Pricing and hedging
Let $d\bar{q}:=dq-qr\,dt$ be the underlying return centralised in the price measure. Using the price measure, the hedge and price conditions are re-expressed as:
00 0 $=\mathbb{E}^{\phi}[(d\bar{q}-\alpha qs\,dt)\exp[-\alpha(P-\delta\cdot d\bar{q})]]$ (54)
$p$ $=\frac{1}{1+r\,dt}\mathbb{E}^{\phi}[\alpha][P-\delta\cdot d\bar{q}]$
@@ III.2 Pricing and hedging
00 0 $=\mathbb{E}^{\phi}[(d\bar{q}-\lambda\alpha qs\,dt)\exp[-\lambda\alpha(P-\delta\cdot d\bar{q})]]$ (55)
$p$ $=\frac{1}{1+r\,dt}\mathbb{E}^{\phi}[\lambda\alpha][P-\delta\cdot d\bar{q}]$
@@
$s$ $=\frac{\displaystyle\mathbb{E}^{\phi}[P]/(1+r\,dt)-\mathbb{V}^{\phi}[dq]^{-1}\mathbb{V}^{\phi}[dq,P]\cdot q}{\displaystyle\mathbb{V}^{\phi}[dq]^{-1}q\,dt\cdot q}+O[\lambda\alpha]$
@@
$p$ $=\frac{1}{1+r\,dt}\left(\mathbb{E}^{\phi}[P]-\frac{1}{2}\lambda\alpha\mathbb{V}^{\phi}[P-\delta\cdot dq]\right)+O[(\lambda\alpha)^{2}]$
@@
$p_{\text{mid}}$ $=\mathbb{E}^{\phi}[P]/(1+r\,dt)$ (57)
$p_{\text{bid-offer}}$ $=\alpha\mathbb{V}^{\phi}[P-\delta\cdot dq]/(1+r\,dt)$
@@
This model is extensive in scope, consistently encompassing arbitrary economic assumptions, predictable and unpredictable funding, complete and incomplete hedge markets, continuous and discontinuous evolution, continuous and discrete hedging, and accounting for the risk appetite of the investor. The studies in figure 7 demonstrate entropic hedging and pricing for call and digital options over a single investment horizon, and show the bid-offer around the mid model that prices the residual risk after hedging.
It is interesting to note that entropic pricing handles the trivial case of an empty underlying market in a sensible way. The calibration and hedge conditions are empty, and the price condition implies that the derivative price is a log-martingale in the expectation measure:
$p=\mathbb{E}[\alpha][P]$ (58)
In the absence of any market information, price reduces to expectation adjusted to match the risk appetite of the investor.
When the underlying market is not empty, the hedge and price conditions may not have simple solutions and are instead solved numerically via the Newton-Raphson scheme. The conditions can be solved explicitly when the prices diffuse continuously, however, and this case is useful to develop intuition. The diffusion of the underlying and derivative prices is continuous in the expectation measure $\mathbb{E}$ when their cumulant generating function satisfies:
$\log{}$ $\mathbb{E}\exp[\beta\,dp+\psi\cdot dq]=$ (59)
$\beta\mathbb{E}[dp]+\psi\cdot\mathbb{E}[dq]$
$+\frac{1}{2}\beta^{2}\mathbb{V}[dp]+\beta\psi\cdot\mathbb{V}[dq,dp]+\frac{1}{2}\psi\cdot\mathbb{V}[dq]\psi+O[dt^{2}]$
with $\mathbb{V}[dp]=O[dt]$, $\mathbb{V}[dq,dp]=O[dt]$ and $\mathbb{V}[dq]=O[dt]$. If the investor re-balances frequently and at negligible cost, the $O[dt]$ corrections to the investment strategy can be ignored, and the bid-offer around the mid price derives exclusively from the post-hedging residual risks.
Moving to the price measure $\mathbb{E}^{\phi}$, the impact of the measure change on the cumulant generating function derives the drift adjustments in Girsanov’s theorem:
$\log{}$ $\mathbb{E}^{\phi}\exp[\beta\,dp+\psi\cdot dq]=$ (60)
$\beta(\mathbb{E}[dp]-\phi\cdot\mathbb{V}[dq,dp])+\psi\cdot(\mathbb{E}[dq]-\mathbb{V}[dq]\phi)$
$+\frac{1}{2}\beta^{2}\mathbb{V}[dp]+\beta\psi\cdot\mathbb{V}[dq,dp]+\frac{1}{2}\psi\cdot\mathbb{V}[dq]\psi+O[dt^{2}]$
Differentiating with respect to $\psi$ then generates:
$\frac{\displaystyle\mathbb{E}^{\phi}[dq\,\exp[\beta\,dp+\psi\cdot dq]]}{\displaystyle\mathbb{E}^{\phi}[\exp[\beta\,dp+\psi\cdot dq]]}=$ (61)
$\qquad\mathbb{E}[dq]+\beta\mathbb{V}[dq,dp]+\mathbb{V}[dq](\psi-\phi)+O[dt^{2}]$
These results are sufficient to solve the conditions. The calibration condition is solved:
$r$ $=\frac{\displaystyle\mathbb{V}[dq]^{-1}\mathbb{E}[dq]\cdot q}{\displaystyle\mathbb{V}[dq]^{-1}q\,dt\cdot q}+O[dt]$ (62)
$\phi$ $=\mathbb{V}[dq]^{-1}(\mathbb{E}[dq]-qr\,dt)+O[dt]$
for the implied funding rate $r$ and the unit optimal portfolio $\phi$, and the hedge and price conditions are solved:
$s$ $=\frac{\displaystyle\mathbb{E}^{\phi}[P]-\mathbb{V}[dq]^{-1}\mathbb{V}[dq,P]\cdot q}{\displaystyle\mathbb{V}[dq]^{-1}q\,dt\cdot q}+O[dt]$
$\delta$ $=\mathbb{V}[dq]^{-1}(\mathbb{V}[dq,P]+qs\,dt)+O[dt]$ (63)
$p$ $=(1-r\,dt)\mathbb{E}^{\phi}[P]-\frac{1}{2}\alpha\mathbb{V}[P-\delta\cdot dq]+O[dt^{2}]$
for the implied funding spread $s$, the hedge portfolio $\delta$ and the initial price $p$.
III.3 Market and model risk
Starting with the expectation measure $\mathbb{E}$, the calibration conditions:
00 0 $=\phi\cdot q$ (64)
00 0 $=\mathbb{E}^{\phi}[dq]-qr\,dt$

### v7

@@ III.1 Calibration
  In the second study, the final prices are modelled as independent binomial variables with possible values (0,1+cdt)(0,1+c\,dt) and probabilities (w,1w)(w,1-w). Examples are plotted for the first underlying with coupon $c_{1}=5\%,10\%,15\%$ and default probability $w_{1}=0.01\%,\ldots,10\%$. The second underlying has coupon $c_{2}=10\%$ and the implied funding rate is plotted as a function of its default probability $w_{2}$. The implied funding rate is capped by the coupons and decreases as the default probability increases.
While there is no explicit solution to the calibration condition, the unit optimal portfolio and implied funding rate can be identified in some scenarios, and more generally are efficiently discovered by iterating the Newton-Raphson scheme. In each of these scenarios, the expectation in the calibration condition is simplified or approximated, which facilitates its solution.
Normal model
When the underlying prices are joint normally distributed in the expectation measure, their cumulant generating function is:
$\log\mathbb{E}\exp[\psi\cdot dq]$ $=$ (68)
$\hskip-56.9055pt\psi\cdot\mathbb{E}[dq]+\frac{1}{2}\psi\cdot\mathbb{V}[dq]\psi$
$\log\mathbb{E}^{\phi}\exp[\psi\cdot dq]$ $=$
$\hskip-56.9055pt\psi\cdot(\mathbb{E}[dq]-\mathbb{V}[dq]\phi)+\frac{1}{2}\psi\cdot\mathbb{V}[dq]\psi$
The calibration condition becomes:
$qr\,dt=\mathbb{E}[dq]-\mathbb{V}[dq]\phi$ (69)
which is solved with the unit optimal portfolio portfolio:
$\phi=\mathbb{V}[dq]^{-1}(\mathbb{E}[dq]-qr\,dt)$ (70)
The implied funding rate is then determined from the self-funding condition:
$r\,dt=\frac{\displaystyle\mathbb{V}[dq]^{-1}q\cdot\mathbb{E}[dq]}{\displaystyle\mathbb{V}[dq]^{-1}q\cdot q}$ (71)
Continuous model
When the underlying prices diffuse continuously in the expectation measure, their cumulant generating function is:
$\log\mathbb{E}\exp[\psi\cdot dq]$ $=$ (72)
$\hskip-56.9055pt\psi\cdot\mathbb{E}[dq]+\frac{1}{2}\psi\cdot\mathbb{V}[dq]\psi+O[dt^{2}]$
$\log\mathbb{E}^{\phi}\exp[\psi\cdot dq]$ $=$
$\hskip-56.9055pt\psi\cdot(\mathbb{E}[dq]-\mathbb{V}[dq]\phi)+\frac{1}{2}\psi\cdot\mathbb{V}[dq]\psi+O[dt^{2}]$
where the first-order cumulants are $O[1]$ or higher, the second-order cumulants are $O[dt]$ and the higher-order cumulants are $O[dt^{2}]$ or higher. The calibration condition becomes:
$qr\,dt=\mathbb{E}[dq]-\mathbb{V}[dq]\phi+O[dt^{2}]$ (73)
which is solved with the unit optimal portfolio portfolio:
$\phi=\mathbb{V}[dq]^{-1}(\mathbb{E}[dq]-qr\,dt)+O[dt]$ (74)
The implied funding rate is then determined from the self-funding condition:
$r\,dt=\frac{\displaystyle\mathbb{V}[dq]^{-1}q\cdot\mathbb{E}[dq]}{\displaystyle\mathbb{V}[dq]^{-1}q\cdot q}+O[dt^{2}]$ (75)
In addition to the assumption that the underlying prices diffuse continuously, this solution is applicable only when the optimal portfolio is rebalanced with sufficient frequency so that the $O[dt]$ corrections to the strategy can be neglected.
Numerical scheme
At each step the Newton-Raphson scheme improves the approximate unit optimal portfolio $\phi$ by the correction $\partial\phi$ which solves the calibration and self-funding conditions to $O[\partial\phi]$. Neglecting $O[\partial\phi^{2}]$ and higher terms, the calibration condition becomes:
$qr\,dt=\mathbb{E}^{\phi}[dq]-\mathbb{V}^{\phi}[dq]\,\partial\phi$ (76)
which is solved with the correction:
$\partial\phi=\mathbb{V}^{\phi}[dq]^{-1}(\mathbb{E}^{\phi}[dq]-qr\,dt)$ (77)
The implied funding rate $r$ is then determined from the self-funding condition:
$r\,dt=\frac{\displaystyle(\phi+\mathbb{V}^{\phi}[dq]^{-1}\mathbb{E}^{\phi}[dq])\cdot q}{\displaystyle\mathbb{V}^{\phi}[dq]^{-1}q\cdot q}$ (78)
The Newton-Raphson scheme is initialised with the choice $\phi=0$ and repeats these steps until the calibration condition is satisfied to the required level of accuracy.
@@
The implied funding rate incorporates contributions from all the underlying returns, weighted by their predictabilities so that the least risky returns dominate. This property is demonstrated in the studies of figure˜8, which simplify the model with the assumption that the final underlying prices are independent. The calibration conditions decouple to generate a set of independent conditions, with the $i$th underlying security satisfying:
@@
These conditions are solved separately to express the components of the unit optimal portfolio $\phi[r]$ as functions of the implied funding rate $r$. The self-funding condition $\phi[r]\cdot q=0$ is then solved for $r$.
In the first study, the final underlying price $Q_{i}$ is modelled as a normal variable with mean $q_{i}(1+\mu_{i}\,dt)$ and variance $q_{i}^{2}\sigma_{i}^{2}\,dt$, where $\mu_{i}$ is the drift and $\sigma_{i}$ is the volatility. The calibration condition derives:
@@
and the self-funding condition then derives:
@@
In the second study, the final underlying price $Q_{i}$ is modelled as a binomial variable with possible values $(0,1+c_{i}\,dt)$ occurring with probabilities $(w_{i},1-w_{i})$, where $c_{i}$ is the coupon and $w_{i}$ is the default probability. The calibration condition derives:
@@
III.2 Hedging and pricing
The price condition determines the initial price $p$ from the final price $P$ of the derivative security via the relationship:
$0=\mathbb{E}^{\phi}[\alpha][dp-\delta\cdot dq]$ (84)
where the investment is funded by the hedge portfolio $\delta$ in the underlying market, chosen to maximise the entropy-adjusted mean return at risk aversion $\alpha$ subject to the self-funding condition $\delta\cdot q=p$.
When $\alpha=0$ the investor neglects the impact of residual risk from the hedging strategy and the derivative price is a discounted martingale in the price measure:
$p=\frac{1}{1+r\,dt}\mathbb{E}^{\phi}[P]$ (85)
Over the discrete schedule $t_{0}<t_{1}<\cdots<t_{n}$ the price condition is compounded to derive:
$p_{i}=\mathbb{E}_{i}^{\phi}\!\left[\prod_{j=i}^{n-1}\frac{1}{1+r_{j}\,dt_{j}}p_{n}\right]$ (86)
The model bears all the hallmarks of risk-neutral theory, including a kernel that performs stochastic discounting and a price measure that is equivalent to the expectation measure, but applies more generally. Entropic risk optimisation uniquely identifies both these ingredients from the optimal strategy in the underlying market, and does not need the additional assumptions of market completeness and the existence of risk-free funding.
When $\alpha\neq 0$ the investor reserves against residual risk from the hedging strategy and the derivative price is a discounted log-martingale in the price measure:
$p=\mathbb{E}^{\phi}[\alpha][P-\delta\cdot dq]$ (87)
Over the discrete schedule $t_{0}<t_{1}<\cdots<t_{n}$ the price condition is compounded to derive:
$p_{i}=\mathbb{E}_{i}^{\phi}[\alpha]\!\left[p_{n}-\sum_{j=i}^{n-1}\delta_{j}\cdot dq_{j}\right]$ (88)
The model adjusts the initial price to protect the hedging strategy from the negative consequences of unhedged risks. If the underlying market is complete for the derivative security, so that the return on the hedged portfolio has zero variance, then no buffer is required and the price simplifies to the risk-neutral solution.
Hedging identifies the price measure $\mathbb{E}^{\phi\delta}:=\mathbb{E}^{(\phi-\alpha\delta)\oplus\alpha}$ equivalent to the price measure $\mathbb{E}^{\phi}$ with Radon-Nikodym weight:
$\frac{\displaystyle d\mathbb{E}^{\phi\delta}}{\displaystyle d\mathbb{E}^{\phi}\hfill}:=\frac{\displaystyle\exp[-\alpha(P-\delta\cdot dq)]}{\displaystyle\mathbb{E}^{\phi}[\exp[-\alpha(P-\delta\cdot dq)]]}$ (89)
where the hedge portfolio $\delta$ and implied funding spread $s$ are solved from the hedge and self-funding conditions:
$q(r+\alpha s)\,dt$ $=\mathbb{E}^{\phi\delta}[dq]$ (90)
$\delta\cdot q$ $=\mathbb{E}^{\phi}[\alpha][P-\delta\cdot dq]$
The implied funding spread accounts for the incremental funding cost or benefit arising from the inclusion of the derivative security. Given the hedge portfolio $\delta$, the initial derivative price $p=\delta\cdot q$ decomposes as the weighted sum of its component initial underlying prices. The hedge portfolio of underlying securities thus simultaneously funds and hedges the derivative security, and is optimally selected for both objectives.
While there is no explicit solution to the hedge condition, the hedge portfolio and implied funding spread can be identified in some scenarios, and more generally are efficiently discovered by iterating the Newton-Raphson scheme. In each of these scenarios, the expectations in the hedge and self-funding conditions are simplified or approximated, which facilitates their solution.
Normal model
When the underlying and derivative prices are joint normally distributed in the expectation measure, their cumulant generating function is:
$\log\mathbb{E}\exp[\beta\,dp+\psi\cdot dq]$ $=$ (91)
$\hskip-85.35826pt\beta\mathbb{E}[dp]+\psi\cdot\mathbb{E}[dq]$
$\hskip-85.35826pt+\frac{1}{2}\beta^{2}\mathbb{V}[dp]+\beta\psi\cdot\mathbb{V}[dq,dp]+\frac{1}{2}\psi\cdot\mathbb{V}[dq]\psi$
$\log\mathbb{E}^{\phi}\exp[\beta\,dp+\psi\cdot dq]$ $=$
$\hskip-85.35826pt\beta(\mathbb{E}[dp]-\phi\cdot\mathbb{V}[dq,dp])+\psi\cdot(\mathbb{E}[dq]-\mathbb{V}[dq]\phi)$
$\hskip-85.35826pt+\frac{1}{2}\beta^{2}\mathbb{V}[dp]+\beta\psi\cdot\mathbb{V}[dq,dp]+\frac{1}{2}\psi\cdot\mathbb{V}[dq]\psi$
With this assumption, the calibration condition is solved with unit optimal portfolio $\phi=\mathbb{V}[dq]^{-1}(\mathbb{E}[dq]-qr\,dt)$ where the implied funding rate $r$ is tuned to the condition $\phi\cdot q=0$. The hedge condition becomes:
$q(r+\alpha s)\,dt=qr\,dt-\alpha\mathbb{V}[dq,P]+\alpha\mathbb{V}[dq]\delta$ (92)
which is solved with the hedge portfolio:
$\delta=\mathbb{V}[dq]^{-1}(\mathbb{V}[dq,P]+qs\,dt)$ (93)
The self-funding condition becomes:
$\delta\cdot q=$ (94)
$\quad\mathbb{E}[P]-\mathbb{V}[dq]^{-1}(\mathbb{E}[dq]-qr\,dt)\cdot\mathbb{V}[dq,P]-\frac{1}{2}\alpha\mathbb{V}[P]$
$\quad-\delta\cdot(qr\,dt-\alpha\mathbb{V}[dq,P])-\frac{1}{2}\alpha\delta\cdot\mathbb{V}[dq]\delta$
which is solved with the implied funding spread:
$s\,dt={}$ $\frac{1}{\alpha}(1+r\,dt)\times$ (95)
$\left(\sqrt{1+2\alpha\frac{\displaystyle\mathbb{E}[\alpha][P-\mathbb{V}[dq]^{-1}\mathbb{V}[dq,P]\cdot Q]}{\displaystyle\mathbb{V}[dq]^{-1}q\cdot q(1+r\,dt)^{2}}}-1\right)$
The initial derivative price is then given by $p=\delta\cdot q$. This solution applies when the underlying and derivative prices are joint normally distributed, and is a useful approximation in near-normal scenarios.
Continuous model
When the underlying and derivative prices diffuse continuously in the expectation measure, their cumulant generating function is:
$\log\mathbb{E}\exp[\beta\,dp+\psi\cdot dq]$ $=$ (96)
$\hskip-85.35826pt\beta\mathbb{E}[dp]+\psi\cdot\mathbb{E}[dq]$
$\hskip-85.35826pt+\frac{1}{2}\beta^{2}\mathbb{V}[dp]+\beta\psi\cdot\mathbb{V}[dq,dp]+\frac{1}{2}\psi\cdot\mathbb{V}[dq]\psi+O[dt^{2}]$
$\log\mathbb{E}^{\phi}\exp[\beta\,dp+\psi\cdot dq]$ $=$
$\hskip-85.35826pt\beta(\mathbb{E}[dp]-\phi\cdot\mathbb{V}[dq,dp])+\psi\cdot(\mathbb{E}[dq]-\mathbb{V}[dq]\phi)$
$\hskip-85.35826pt+\frac{1}{2}\beta^{2}\mathbb{V}[dp]+\beta\psi\cdot\mathbb{V}[dq,dp]+\frac{1}{2}\psi\cdot\mathbb{V}[dq]\psi+O[dt^{2}]$
where the first-order cumulants are $O[1]$ or higher, the second-order cumulants are $O[dt]$ and the higher-order cumulants are $O[dt^{2}]$ or higher. With this assumption, the calibration condition is solved with unit optimal portfolio $\phi=\mathbb{V}[dq]^{-1}(\mathbb{E}[dq]-qr\,dt)+O[dt]$ where the implied funding rate $r$ is tuned to the condition $\phi\cdot q=0$. The hedge condition becomes:
$q(r+\alpha s)\,dt=qr\,dt-\alpha\mathbb{V}[dq,P]+\alpha\mathbb{V}[dq]\delta+O[dt^{2}]$ (97)
which is solved with the hedge portfolio:
$\delta=\mathbb{V}[dq]^{-1}(\mathbb{V}[dq,P]+qs\,dt)+O[dt]$ (98)
The self-funding condition becomes:
$\delta\cdot q=\mathbb{E}[P]-\mathbb{V}[dq]^{-1}\mathbb{E}[dq]\cdot\mathbb{V}[dq,P]+O[dt]$ (99)
which is solved with the implied funding spread:
$s\,dt=\frac{\displaystyle\mathbb{E}[P]-\mathbb{V}[dq]^{-1}\mathbb{V}[dq,P]\cdot\mathbb{E}[Q]}{\displaystyle\mathbb{V}[dq]^{-1}q\cdot q}+O[dt^{2}]$ (100)
The initial derivative price is then given by:
$p={}$ $\frac{1}{1+r\,dt}\mathbb{E}^{\phi}[P]$ (101)
$-\frac{1}{2}\alpha\mathbb{V}[P-\mathbb{V}[dq]^{-1}\mathbb{V}[dq,P]\cdot dq]+O[dt^{2}]$
In addition to the assumption that the underlying and derivative prices diffuse continuously, this solution is applicable only when the hedge portfolio is rebalanced with sufficient frequency so that the $O[dt]$ corrections to the strategy can be neglected.
Numerical scheme
At each step the Newton-Raphson scheme improves the approximate hedge portfolio $\delta$ by the correction $\partial\delta$ which solves the hedge and self-funding conditions to $O[\partial\delta]$. Neglecting $O[\partial\delta^{2}]$ and higher terms, the hedge condition becomes:
$q(r+\alpha s)\,dt=\mathbb{E}^{\phi\delta}[dq]+\alpha\mathbb{V}^{\phi\delta}[dq]\,\partial\delta$ (102)
which is solved with the correction:
$\partial\delta=-\frac{1}{\alpha}\mathbb{V}^{\phi\delta}[dq]^{-1}(\mathbb{E}^{\phi\delta}[dq]-q(r+\alpha s)\,dt)$ (103)
Neglecting $O[\partial\delta^{2}]$ and higher terms, the self-funding condition becomes:
$(\delta+\partial\delta)\cdot q=\mathbb{E}^{\phi}[\alpha][P-\delta\cdot dq]-\partial\delta\cdot\mathbb{E}^{\phi\delta}[dq]$ (104)
which is solved with the implied funding spread:
$(r+\alpha s)\,dt={}$ (105)
$\hskip-28.45274pt\frac{\displaystyle\alpha\mathbb{E}^{\phi}[\alpha][P-\delta\cdot Q]+\mathbb{V}^{\phi\delta}[dq]^{-1}\mathbb{E}^{\phi\delta}[dq]\cdot\mathbb{E}^{\phi\delta}[Q]}{\displaystyle\mathbb{V}^{\phi\delta}[dq]^{-1}q\cdot\mathbb{E}^{\phi\delta}[Q]}$
The Newton-Raphson scheme is initialised with the choice $\delta=0$ and repeats these steps until the hedge condition is satisfied to the required level of accuracy. The initial derivative price is then given by $p=\delta\cdot q$.
@@ III.2 Hedging and pricing
Using the price measure, the hedge and price conditions are rewritten as:
$q(r+\alpha s)\,dt$ $=\frac{\displaystyle\mathbb{E}^{\phi}[dq\,\exp[-\alpha(P-\delta\cdot dq)]]}{\displaystyle\mathbb{E}^{\phi}[\exp[-\alpha(P-\delta\cdot dq)]]}$ (106)
$p$ $=\mathbb{E}^{\phi}[\alpha][P-\delta\cdot dq]$
@@
$q(r+\lambda\alpha s)\,dt$ $=\frac{\displaystyle\mathbb{E}^{\phi}[dq\,\exp[-\lambda\alpha(P-\delta\cdot dq)]]}{\displaystyle\mathbb{E}^{\phi}[\exp[-\lambda\alpha(P-\delta\cdot dq)]]}$ (107)
$p$ $=\mathbb{E}^{\phi}[\lambda\alpha][P-\delta\cdot dq]$
@@
$s$ $=\frac{\displaystyle\mathbb{E}^{\phi}[P]/(1+r\,dt)-\mathbb{V}^{\phi}[dq]^{-1}\mathbb{V}^{\phi}[dq,P]\cdot q}{\displaystyle\mathbb{V}^{\phi}[dq]^{-1}q\cdot q\,dt}+O[\lambda\alpha]$
@@
$p$ $=\frac{1}{1+r\,dt}\left(\mathbb{E}^{\phi}[P]-\frac{1}{2}\alpha\mathbb{V}^{\phi}[P-\delta\cdot dq]\right)+O[(\lambda\alpha)^{2}]$
@@
$p_{0}$ $=\frac{1}{1+r\,dt}\mathbb{E}^{\phi}[P]$ (109)
$p_{\pm}$ $=\frac{\alpha}{1+r\,dt}\mathbb{V}^{\phi}[P-\delta\cdot dq]$
@@
This model is extensive in scope, consistently encompassing arbitrary economic assumptions, predictable and unpredictable funding, complete and incomplete hedge markets, continuous and discontinuous evolution, continuous and discrete hedging, and accounting for the risk appetite of the investor. The studies in figure˜9 demonstrate entropic hedging and pricing for call and digital options over a single investment horizon, and show the bid-offer around the mid model that prices the residual risk after hedging.
IV Market and model risk
Derivative price models support trading activity, and their risks manifest as poor decision making in the trade lifecycle. The model must provide entry and exit prices that have a reasonable chance of transacting, and a funding and hedging strategy that has a reasonable chance of retaining the initial premium for the trade. The market and model risks in these objectives are distinguished.
Market risk
An optimal funding and hedging strategy is derived on the assumption of the expectation measure. Residual profit and loss is generated if this strategy is not faithfully executed.
Model risk
If the expectation measure does not accurately capture realised market dynamics over the trade lifecycle, then the strategy is suboptimal and will also generate residual profit and loss.
Analysis of these risks considers two expectation measures: the before measure $\mathbb{E}$ assumed initially as the basis for decision making and used to quantify the predicted performance of the strategy; and the after measure $\bar{\mathbb{E}}$ fitted to realised market dynamics and used to quantify the actual performance of the strategy.
Starting with the before expectation measure $\mathbb{E}$, the calibration conditions:
$0$ $=\phi\cdot q$ (110)
$0$ $=\mathbb{E}^{\phi}[dq]-qr\,dt$